OpenStudy (anonymous):

find integral[ x/(1+sin(x))

OpenStudy (zarkon):

I would try the Weierstrass substitution

OpenStudy (anonymous):

what about the "x" in the numerator ?

OpenStudy (zarkon):

\[x=\frac{\tan^{-1}(t)}{2}\] it can be done but it is really nasty you could also try multiplying top and bottom by \[1-\sin(x)\] and write in terms of sec(x) and tan(x)

OpenStudy (anonymous):

ya i tried but the problem always is the "x" in the numerator how to get rid of it ?

OpenStudy (anonymous):

well changing an algebric part to an inverse function not advisory in integration !

OpenStudy (zarkon):

if you use my second suggestion you can integrate by parts to get rid of the x

OpenStudy (anonymous):

following that i get an expression

OpenStudy (anonymous):

\[\[\int\limits\limits_{}^{}\sec ^{-1}t dt\]\]

OpenStudy (zarkon):

\[\int\frac{x}{1+\sin(x)}dx\] \[\int\frac{x}{1+\sin(x)}\frac{1-\sin(x)}{1-\sin(x)}dx\] \[\int\frac{x(1-\sin(x))}{1-\sin^2(x)}dx\] \[\int\frac{x(1-\sin(x))}{\cos^2(x)}dx\] \[\int\frac{x-x\sin(x)}{\cos^2(x)}dx\] \[\int x\cdot\sec^2(x)-x\cdot\sec(x)\tan(x)dx\]